On January 1, 2021, HGC Camera Store adopted the dollar-value LIFO retail inventory method. Inventory transactions at both cost and retail, and cost indexes for 2021 and 2022 are as follows: 2021 2022 Cost Retail Cost Retail Beginning inventory $ 57,000 $ 76,000 Net purchases 109,160 126,000 $ 115,150 $ 132,400 Freight-in 3,800 4,300 Net markups 19,000 11,600 Net markdowns 3,800 4,000 Net sales to customers 124,460 120,440 Sales to employees (net of 10% discount) 4,500 5,850 Price Index: January 1, 2021 1.00 December 31, 2021 1.07 December 31, 2022 1.12
Answer #1

Dollar value LIFO method

For 2021

Cost Retail
Beginning inventory (cost to retail percentage 0.75) 57,000 76,000
Net purchases 109,160 126,000
Freight-in 3,800
Net markups 19,000
(-) net mark downs (3,800)

Goods available for sale (excluding beginning inventory)

{Cost to retail percentage ( 112,960÷141,200)= 0.80}

112,960 141,200
Goods available for sale 169,960   217,200
(-) Net sales ( 124,460+5000) including sales to employees discount (129,460)
Ending inventory at retail 87,740
Ending inventory at cost (62,136)
Estimated cost of goods sold 107,824

Employee discount 4,500

5000-4,500= 500 discount .90

Calculation of ending inventory at cost

Ending inventory at year end retail prices

Step-1 ending inventory at base price step-2 inventory layers at base year price

step -3

Inventory layers converted to cost

87,740 80,000(87,740÷1.07) 76000 57000(76,000×.75)
6000 5,136(6,000×1.07×.8)
62,136

For 2022

Cost Retail
Beginning inventory (cost to retail percentage=.70) 62,136 87,740
Net purchases 115,150 132,400
Freight-in 4,300
Net markups 11,600
(-) net markdowns (4,000)

Goods available for sale (excluding beginning inventory)

{Cost retail percentage (119,450÷140,000)= .85}

119,450 140,000
Goods available for sale( including beginning inventory) 181,586 227,740
(-) Net sales(120,440+6,500) including sales to employees discount (126,940)
Ending inventory at retail 100,800
Ending inventory at cost (63,442)
Estimated cost of goods sold 118,144

Employee discount (10%) 5,850

6,500-650=5,850 employee discount .90

Calculation of ending inventory at cost

Ending inventory at retail price step -1 ending inventory at base price step -2 Inventory layers at base year price step -3 Inventory layers converted to cost
100,800 90,000(100,800÷1.12) 87,740 61,418(87,740×.70)
2,260 2024(2,260×1.12×.85)
63,442

____×____

Comment if you have any doubt

Keep supporting.